OpenStudy (anonymous):

Factor: c^2 - d^2 A. (c + d)(c - d) B. (c + d)(c + d) C. (c + 2d)(c - d)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

a

OpenStudy (anonymous):

If you are unsure do the multiplication: \[(c+d)(c-d)=c^2-cd+cd-d^2 = c^2-d^2\]

OpenStudy (anonymous):

The "difference of two squares" rule is worth remembering, at least until factoring second-degree polynomials becomes second-nature: \[\Large {a^2} - {b^2} = (a + b)(a - b)\]